# How do Pre-Trained Models Support Software Engineering? An Empirical Study in Hugging Face

### v1

@@ III-B Research Method
III-B1 Task Identification
To identify SE tasks applicable to the five stages of the SDLC, we built on prior literature and extended it to form a broader taxonomy. We based our approach on Hou et al. [16], which identifies 88 SE tasks. As a preliminary adjustment, we excluded the task agile story point estimation from the software development activity, and we renamed this activity to software implementation for consistency with other studies, such as [11].
Building upon this foundation, we extended the original list of tasks along two dimensions. First, we broadened the scope of SE activities by refining and expanding their definitions. For example, we explicitly considered software architecture tasks as part of the software design activity. Second, we enriched the set of tasks within each activity by incorporating additional tasks drawn from the Software Engineering Body of Knowledge (SWEBOK) [22], resulting in a total of 147 tasks. Specifically, we used the following SWEBOK chapters: requirements engineering (Chapter 1), software design (Chapters 2 and 3), software implementation (Chapter 4), software quality assurance (Chapters 5, 12, and 13), and software maintenance (Chapter 7).
III-B2 Data Collection
We accessed the HF API [23] to retrieve all PTMs available as of March 24, 2025. For each identifier, we tried to retrieve (i) its model card description, a markdown file documenting the model’s characteristics, intended uses, and evaluation [24]; (ii) the associated card metadata, specified in a YAML block (e.g., license, tags, language) [25]; and (iii) the abstract of a linked arXiv paper [26], taking advantage of the cross-platform linking between HF and arXiv [27]. Any of these components may be missing for a given PTM. At that time, there were 1,533,973 PTMs.
III-B3 Data Preparation
To ensure consistency in downstream analysis, we preprocessed the identified tasks and the collected data by applying tokenization, lowercasing, and lemmatization. Tokenization splits the text into smaller units, such as words, which simplifies the subsequent analysis, while lowercasing eliminates inconsistencies arising from variations in capitalization. Additionally, lemmatization reduces words to their base forms, allowing different inflections to be treated as a single entity. Together, these steps standardize the textual descriptions, ensuring greater accuracy in classification.
III-B4 Data Search
@@
III-B5 Data Filtering
We applied several techniques to refine the available resources and ensure that only the most relevant and accurate ones were included in our classification.
5.1) Outlier Detection: We examined the number of PTMs associated with each task to identify those that significantly deviated from the others. We found that debugging, logging, and coding had notably higher occurrences, together accounting for 89.01% of task occurrences. For debugging, some entries contained debug:None in the metadata, which triggered a match in the search, though it should not have. Similarly, for logging, the lemma “log” often appeared in code snippets or reporting tables, such as those involving logging functions (e.g., import logging) or entries like train log, which were not relevant to the task. For coding, we considered a model to be for coding only if it included the tag “code” in the metadata. To address this, we revisited the raw data for these PTMs without lemmatizing, as we wanted to understand the context in which the lemmas appeared. We checked for any occurrences that did not meet our exclusion criteria (Table III), resulting in a refined set of 14,962 PTMs.
TABLE III: Exclusion criteria for the outlier detection filtering.
@@ III-B5 Data Filtering
5.2) Unique Card Detection: During preliminary analysis, we observed that many model cards exhibit high textual similarity. For instance, the model card of deepseek-ai/DeepSeek-V3-Base[28] is nearly identical to those of other PTMs on the platform, such as deepseek-ai/DeepSeek-V3[29] and Conexis/DeepSeek-R1-NextN-Channel-INT8 [30]. To address this observation, we applied a duplicate detection step using term frequency - inverse document frequency and cosine similarity. We set a 0.99 similarity threshold to capture near-identical cards while allowing for minimal variations, such as the model name. This resulted in 7,797 unique PTMs, with the full list and their high-similarity matches available in the replication package [21].
5.3) Prompt Pilot Test: As part of the filtering stage, we employed an LLM (Gemini 2.0 Flash, selected due to its strong benchmark performance [31]) to assess the relevance of PTMs to SE activities following the PRIMES framework proposed by De Martino et al. [32]. We conducted five pilot tests (one per SE activity) to validate the LLM. For this reason, we selected five random subsets and ensured representation across all their tasks. The sample sizes were determined using a sample size calculator, applying a 95% confidence level and a 5% margin of error to ensure statistical validity[33]. Table IV presents the unique PTMs per SE activity and the corresponding pilot sample sizes. Note that a given PTM can be associated with more than one SE activity.
TABLE IV: PTMs per SE activity and pilot sample sizes.
@@
The first author manually annotated all the pilot samples for each SE activity. In addition to providing a binary judgment (yes/no) on whether a PTM was relevant to a given SE activity, the annotator also documented the underlying reasoning behind the decision. To ensure reliability, 10% of the pilot samples were independently annotated by the second and last authors, both knowledgeable in the domain, drawing on established intercoder reliability practices [34]. To further consolidate our annotation guidelines and understanding, the three annotators participated in a joint discussion session focused on the cases where disagreement had occurred. This discussion revealed that some PTMs are inherently challenging to classify due to their broad or ambiguous descriptions. For example, mradermacher/Regex-AI-Llama-3.2-1B-GGUF[35] was initially assigned to software implementation based on the “code generation” tag included by the author in the card metadata. While the description primarily refers to RegEx generation, this tag was an explicit link to the software implementation activity, and we ultimately retained the label. For the 10% of the pilot samples, we applied majority voting to resolve any discrepancies. The resulting decisions, combined with the initial annotations from the first author, are referred to as human annotations.
We then prompted the LLM to provide both a binary judgment and a rationale for each PTM in the pilot set. We used the default temperature setting of the model to balance determinism and variability in responses, ensuring consistent yet informative outputs. The model’s performance was evaluated by comparing its responses with the human annotations using Cohen’s kappa [36]. This process was repeated iteratively, refining the prompt until we reached almost perfect agreement, as seen in Table V. This process was conducted for each of the five SE activities.
TABLE V: Cohen’s kappa for Humans vs LLM agreements.
@@
5.4) LLM Generalization Test: To evaluate the LLM’s ability to generalize beyond the initial piloting subset and avoid prompt overfitting, we evaluated it on a separate, manually labeled evaluation subset that was not involved in any piloting activities. The full subset was annotated by the first author, and 10% of this subset was independently labeled. We then computed Cohen’s kappa to measure agreement between the human labels and the LLM’s responses on this subset, observing an almost perfect level of agreement ($k=0.88$).
IV Results
In this section, we present the findings for each of the RQs outlined in Section III-A.
IV-A What challenges arise when automatically classifying SE PTMs in HF? (RQ1)
IV-A1 What is the current documentation status of HF SE PTMs?
As summarized in Table VI, 68.52% of the PTMs have a non-empty model card, and particularly 7.32% mention an SE task. This aligns with prior findings that model information can be missing or inaccurate due to the self-reporting nature of model metrics [37].
TABLE VI: Current status of HF PTMs: cards and abstracts.
@@ IV-A1 What is the current documentation status of HF SE PTMs?
arXiv Not available or no SE tasks 1,531,643 99.85%
abstract Available, with SE tasks 2,330 0.15%
@@
IV-A2 What are the key findings regarding the correct classification of SE PTMs?
The application of our classification pipeline revealed that a significant number of model cards are duplicated. This is primarily due to two reasons: either they correspond to different variants of the same model (e.g., optimized or finetuned versions), or users re-upload PTMs under their own namespace without making significant modifications to the card content. These duplications reflect the existence of an ecosystem in which PTMs depend on and interact with each other [17]. For example, we observed that sparsh35/Meta-Llama-3.1-8B-Instruct [38] has 290 highly similar counterparts, such as triplee/torchtune_8B_full_finetuned_llama3.
1_millfield_241223_meta_before_user_epoch2[39]
. Similarly, the PTM dducnguyen/newdream-sdxl-
20[40]
is associated with 715 near-duplicate PTMs. To mitigate this, we implemented a duplicate detection strategy whereby the earliest published model card is treated as the original, and all others with high textual similarity are labeled as similar variants of that original.
In addition, during the pilot manual annotation phase, we detected that some tasks can be interpreted in different contexts. These ambiguities were addressed using the LLM to ensure consistency in the classification process. Specifically, the LLM was prompted with explicit inclusion and exclusion criteria for each SE activity, allowing it to disambiguate cases based on contextual clues. Notable examples include:
@@ IV-A2 What are the key findings regarding the correct classification of SE PTMs?
IV-A3 How stable is this classification information over time?
Figure 2 shows a growth in the creation of SE PTMs since 2020, with quarterly data revealing periods of accelerated development. Notably, the number of new PTMs developed for software implementation rose from 28 in 2023 Q1 to 142 in 2023 Q2. This upward trend intensified during the last year (from 2024 Q1 to 2024 Q4), peaking at 444 new PTMs in 2024 Q2. In comparison, the current number of PTMs related to software implementation is more than 20 times higher than that of the next most represented SE activity, software maintenance. In parallel, we observe a steady rise in the number of PTMs associated with requirements engineering, software maintenance, software quality assurance, and software design, indicating a broader diversification of SE tasks being addressed.
Figure 2: Number of SE PTMs created quarterly since 2020.
Finding 1.4: Since 2020, SE PTMs have grown steadily, with sharp acceleration observed from 2023 onward. Finding 1.5: The number of PTMs addressing SE activities continues to grow, with software implementation maintaining a strong lead.
IV-B How are PTMs utilized for automating SE tasks (RQ2)
IV-B1 What SE activities and tasks are covered by PTMs?
Figure 3: PTMs associated with each SE task and SE activity.
@@ IV-B1 What SE activities and tasks are covered by PTMs?
Figure 3 shows the distribution of PTMs across SE tasks, with each bar color-coded to reflect its corresponding SE activity. A clear pattern emerges: the most common SE tasks are code generation and coding. Notably, PTMs designed for code generation outnumber those for general coding support by more than a twofold margin. From an activity-level perspective, the data reveal that software implementation is the most dominant SE activity in terms of PTM coverage. In contrast, early-stage activities such as requirements engineering and software design are less represented, with only a small fraction of PTMs targeting them. Later phases of the SDLC, including software quality assurance and maintenance, demonstrate moderate PTM adoption. Tasks like cybersecurity and sentiment analysis are relatively well-supported, indicating a recognition of the value of automation in ensuring software security.
@@
Finding 2.1: Code generation is the most covered SE task. Finding 2.2: Software implementation-related PTMs dominate, while early stages of the SDLC are underrepresented.
IV-B2 How are ML tasks related to SE activities?
The relationship between SE activities and ML tasks is illustrated in Figure 4, with the flow indicating the number of PTMs. The figure highlights the prominence of the ML task text generation, which is associated with the highest number of PTMs, reflecting its relevance in supporting code-related tasks. From the SE perspective, this task is most commonly linked to software implementation. However, it also extends across the remaining four SE activities, indicating its versatility.
Figure 4: Association of SE activities with ML tasks.
ML support is unevenly distributed across the SDLC. Early phases (requirements engineering and software design) and late phases (software maintenance) demonstrate connections to diverse ML tasks, while middle phases like software design show limited ML task integration, being connected primarily to text generation. In particular, software design is associated with a single ML task (text generation) while the other SE activities are supported by multiple ML tasks. Software quality assurance presents an interesting case with connections to several ML tasks, but notably fewer than software implementation or software maintenance. This highlights a potential growth area for ML applications in testing and validation tasks. Regarding software maintenance, it spans across various ML tasks, including text generation, text classification, and text2text generation. These ML tasks are useful in the context of SE, such as code review tasks.
Finding 2.3: The ML task text generation is the most common among SE-related PTMs. Finding 2.4: Unlike other SE activities, software design is related to a single ML task: text generation.
IV-B3 How are cross-activity PTMs applied across multiple SE activities?
PTMs can be related to more than one SE task, potentially spanning different SE activities (i.e., a PTM can be simultaneously classified into multiple SE activities).

### v2

@@ 3.2 Research Method
3.2.1 Task Identification
To identify SE tasks applicable to the five stages of the SDLC, we built on prior literature and extended it to form a broader taxonomy. We based our approach on 10.1145/3695988, which identifies 88 SE tasks. As a preliminary adjustment and after discussion among the authors, we excluded the task agile story point estimation from the software development activity, and we renamed this activity to software implementation for consistency with other authors, such as sommerville2015.
Building upon this foundation, we extended the original list of tasks along two dimensions. First, we broadened the scope of SE activities by refining and expanding their definitions. For example, we explicitly considered software architecture tasks as part of the software design activity. Second, we enriched the set of tasks within each activity by incorporating additional tasks drawn from the Software Engineering Body of Knowledge (SWEBOK) washizaki2024swebok. Specifically, we considered key concepts and topics described in relevant chapters: requirements engineering (Chapter 1), software design (Chapters 2 and 3), software implementation (Chapter 4), software quality assurance (Chapters 5, 12, and 13), and software maintenance (Chapter 7). The resulting taxonomy, refined through discussion among the authors to ensure coverage and non-overlap, comprises a total of 147 tasks.
3.2.2 Data Collection
We accessed the HF API huggingfaceHuggingFace to retrieve all PTMs available as of March 24, 2025. For each identifier, we tried to retrieve (i) its model card description, a markdown file documenting the model’s characteristics, intended uses, and evaluation mitchell2019model; (ii) the associated card metadata, specified in a YAML block (e.g., license, tags, language) huggingfaceModelCardsMetadata, which has been shown to help users find appropriate PTMs suryani2025model; and (iii) the abstract of a linked arXiv paper huggingfaceModelCardsPaper, taking advantage of the cross-platform linking between HF and arXiv suryani2024exploration. Any of these components may be missing for a given PTM. At that time, there were 1,533,973 PTMs available in HF.
3.2.3 Data Processing
The third phase of the methodology involves four sequential steps to process the collected data, ranging from text normalization (used to standardize matches during task detection) to identifying outliers and unique instances, ensuring that no PTMs are duplicated or counted twice.
3.1) Text Normalization
We preprocessed the list of SE tasks and the collected data to ensure accurate SE task detection across heterogeneous textual sources. Model cards, metadata, and abstracts in HF exhibit diverse writing styles and formats, as terms may appear in uppercase or inflected forms (e.g., “Code Summarization”, “code summarisation” or “code summarizing”). To handle this variability, we applied tokenization, lowercasing, and lemmatization. Together, these steps standardize the textual descriptions, ensuring greater catalogue accuracy.
3.2) Task Search
@@ 3.2.3 Data Processing
3.3) Outlier Detection
We applied several techniques to refine the available resources and ensure that only the most relevant and accurate ones were included in our catalogue. Our primary focus was on detecting positive outliers, i.e., SE tasks whose frequency was abnormally high compared to the rest, as these cases typically signal false positives caused by lexical ambiguities rather than actual SE relevance. We found that debugging, logging, and coding had notably higher occurrences, together accounting for 89.01% of task occurrences. For debugging, some entries contained debug:None in the metadata, which triggered a match in the search, though it should not have. Similarly, for logging, the lemma “log” often appeared in code snippets or reporting tables, such as those involving logging functions (e.g., import logging) or entries like train log, which were not relevant to the task. For coding, we considered a model to be for coding only if it included the tag “code” in the metadata. To ensure robustness of the filtering, we revisited the raw data for these PTMs without lemmatizing, as we wanted to understand the context in which the lemmas appeared, applying the exclusion criteria in Table 3. We then checked for any occurrences that did not meet our exclusion criteria, resulting in a refined set of 14,962 PTMs.
Table 3: Exclusion criteria for the outlier detection filtering
@@
3.4) Unique Card Detection
During preliminary analysis, we observed that many model cards exhibit high textual similarity. For instance, the model card of deepseek-ai/DeepSeek-V3-Base huggingfaceDeepseekaiDeepSeekV3BaseHugging is nearly identical to those of other PTMs on the platform, such as deepseek-ai/DeepSeek-V3 huggingfaceDeepseekaiDeepSeekV3Hugging and Conexis/DeepSeek-R1-NextN-Channel-INT8 huggingfaceConexisDeepSeekR1NextNChannelINT8Hugging. In particular, duplicate and near-duplicate PTMs often emerge because users re-upload base models with little or no modification horwitz2025charting. To address this observation, we applied a duplicate detection step using term frequency - inverse document frequency and cosine similarity. We set a 0.99 similarity threshold, which is high enough to detect near-identical cards while allowing true distinct versions to remain separate because their documentation differs sufficiently to fall below this threshold. Without this step, we risk double-counting PTMs that have minimal documentation differences but are otherwise the same model. This procedure resulted in 7,797 unique PTMs, with the full list and their high-similarity matches available in the replication package replication_package.
3.2.4 LLM-Based Validation
In the fourth step of the methodology, we used an LLM to validate the cataloguing and ensure that we only keep SE-relevant resources. To guarantee objectivity, we conducted a two-phase validation procedure, as illustrated in Figure 2. All prompt iterations and manually annotated subsets are available in the replication package replication_package.
Figure 2: Process of using an LLM to validate the catalogue
4.1) Prompts Refinement
We employed Gemini 2.0 Flash (selected due to its strong benchmark performance rahman2025comparative) to assess the relevance of PTMs to SE activities following the PRIMES framework proposed by de2024experiences.
We conducted five pilot tests (one per SE activity) to validate the LLM. For this reason, we selected five random subsets and ensured representation across all their tasks. The sample sizes were determined using a sample size calculator, applying a 95% confidence level and a 5% margin of error to ensure statistical validity qualtricsSampleSize. Table 4 presents the unique PTMs per SE activity and the corresponding pilot sample sizes. Note that a given PTM can be associated with more than one SE activity.
The first author manually annotated all pilot samples for each SE activity. In addition to providing a binary judgment (yes/no) on whether a PTM was relevant to a given SE activity, the annotator also documented the underlying reasoning behind the decision. To ensure reliability, 10% of the pilot samples were independently annotated by the second and fourth authors, both knowledgeable in the domain, drawing on established intercoder reliability (ICR) practices in qualitative research o2020intercoder. To consolidate our annotation guidelines and understanding, the three annotators participated in a joint discussion session focused on the cases where disagreement had occurred. This discussion revealed that some PTMs are inherently challenging to classify due to their broad or ambiguous descriptions. For example, mradermacher/Regex-AI-Llama-3.2-1B-GGUF huggingfaceMradermacherRegexAILlama321BGGUFHugging was initially assigned to software implementation based on the “code generation” tag included by the author in the card metadata. While the description primarily refers to RegEx generation, this tag was an explicit link to the software implementation activity, and we ultimately retained the label. For the 10% of the pilot samples, we applied majority voting to resolve any discrepancies. The resulting decisions, combined with the initial annotations from the first author, are referred to as the ground truth.
Table 4: PTMs per SE activity and pilot sample sizes
@@ 3.2.4 LLM-Based Validation
We then prompted the LLM in a zero-shot setting to provide both a binary judgment and a rationale for each PTM in the pilot set. We used the default temperature setting of the model to balance determinism and variability in responses, ensuring consistent yet informative outputs.
Prompt Structure The prompt followed a structured format comprising: 1. Objective: Assess the PTM’s applicability to the target activity. 2. Background context: Provide a brief overview of the SDLC and the specific SE activity being evaluated. 3. Task description: Instructions on how to review and assess the PTM for the target SE activity. 4. Criteria: Explicit rules for determining relevance to the target SE activity. 5. Example output: Format for binary judgment and rationale.
The model’s performance was evaluated by comparing its responses with the human annotations using Cohen’s kappa sim2005kappa. This process was repeated iteratively, refining the prompt until we reached almost perfect agreement, as seen in Table 5. This process was conducted for each of the five SE activities. In total, 5 prompt refinements were required for requirements engineering, 12 for software design, 4 for software implementation, 9 for software quality assurance, and 8 for software maintenance. The refinements mainly focused on providing concrete examples of SE tasks for each activity, explicitly instructing the use of the criteria list when making judgments, and avoiding superficial or ambiguous associations with the activity based on the information provided.
Table 5: Cohen’s kappa for Humans vs LLM agreements
@@
4.2) Generalization Test
To evaluate the LLM’s ability to generalize beyond the initial piloting subset and avoid prompt overfitting, we evaluated it on a separate, manually labeled evaluation subset that was not involved in any piloting activities. This set consisted of 366 randomly sampled PTMs across all SE activities, a statistically significant size for our population. The ground truth for this subset was created using the same rigorous process as before: the full subset was annotated by the first author, and 10% of this subset was independently labeled. Then, we instructed the LLM with the five refined prompts and computed Cohen’s kappa to measure agreement between human decisions and the LLM’s responses on this subset, observing an almost perfect level of agreement ($k=0.88$). This result demonstrates a high degree of generalizability and provides sufficient confidence to use the LLM to label the remaining 6,085 samples.
@@
3.2.5 Data Analysis
The final step of the methodology involves analyzing all the data gathered throughout the cataloguing process, from the initial data dump from HF to the final set of 2,205 SE PTMs.
RQ1 Analysis
RQ1 focuses on identifying the main challenges in automatically cataloguing SE PTMs in HF. To address RQ1.1, we relied on the original set of PTMs: the data collected from HF consisting on 1,533,973 PTMs. Within this data, we examined the completeness and quality of its documentation. We computed descriptive statistics on the availability of model cards and associated arXiv abstracts, and whether they included any SE task match.
To study contextual misclassifications (RQ1.2), we analyzed the frequency distribution of SE task tokens across all PTMs and applied outlier detection to isolate abnormally frequent lemmas. The contextual meaning of each outlier was manually reviewed to determine whether occurrences were relevant or false positives. This combination of quantitative and qualitative analysis provided insights into the documentation quality and ambiguity of SE PTM cataloguing.
RQ2 Analysis
For RQ2, we examined the utilization of PTMs across SE activities, the extent to which PTMs span multiple activities, and their alignment with underlying ML tasks. Each PTM was associated with one or more SE tasks from the validated taxonomy of 2,205 PTMs, enabling the characterization of their distribution throughout the SDLC (RQ2.1).
The cross-activity analysis was conducted by examining co-occurrence patterns between activities to identify PTMs associated with multiple ones (RQ2.2). These patterns were represented as a co-occurrence matrix that allowed examination of relationships among activities without assuming exclusivity.
Furthermore, to study the connection between SE activities and ML tasks, we exploited the task-based organization currently used in HF, which categorizes models according to their underlying ML task types, and computed their joint distributions with the SE catalogue (RQ2.3).
RQ3 Analysis
The analyses under RQ3 characterized SE PTMs through temporal, structural, and legal dimensions of SE PTMs. For temporal evolution (RQ3.1), we parsed the creation date field of each PTM, normalized timestamps into quarterly intervals, and computed cumulative growth curves to describe the evolution of SE-related models over time.
To analyze base-model reuse (RQ3.2), we extracted the base_model attribute, introduced by HF in November 2023, and aggregated PTMs by base model and creation date to assess dependency patterns among variants of the same base model (e.g., fine-tuned or derivative models).
For benchmark characterization (RQ3.3), we parsed the model-index section of each card to identify evaluation results. Since naming conventions across model cards are heterogeneous, we applied a rule-based normalization to standardize benchmark and metric names (e.g., unifying benchmark variants such as MMLU and Measuring Massive Multitask Language Understanding, or metrics such as accuracy_norm and normalized_accuracy). This analysis is motivated by the growing interest in SE benchmarks in recent years: for SE-focused LLMs, the number of benchmarks increased from 30 in 2022 to 81 in 2024 hu2025assessing.
To analyze training datasets (RQ3.4), we processed the datasets field in the cards to find patterns in dataset reuse. We mapped dataset usage trends and their alignment with SE activities.
Finally, for licensing (RQ3.5), we extracted license information to assess reuse conditions. Licensing considerations have become increasingly important in the development of PTMs, as recent studies highlight the need for license awareness to mitigate copyright risks lo2023trustworthy. Accordingly, we checked how permissive, copyleft, and restrictive licenses are used per SE activities.
4 Results
In this section, we present the findings from our study, organized according to the three RQs introduced in Section 3.1.
4.1 What challenges arise when automatically cataloguing SE PTMs in HF? (RQ1)
To address RQ1, we examine the main challenges encountered when automatically cataloguing SE PTMs in HF. We first analyze the current documentation status of SE PTMs to assess the completeness and reliability of available model information. Next, we discuss the key findings that emerged during the classification process.
4.1.1 What is the current documentation status of SE PTMs?
As summarized in Table 6, 68.52% of the PTMs have a non-empty model card, and particularly 7.32% mention an SE task. This aligns with prior findings that model information can be missing or inaccurate due to the self-reporting nature of model metrics jiang2023empirical.
Table 6: Availability and content of model cards and arXiv abstracts for PTMs in HF
@@ 4.1.1 What is the current documentation status of SE PTMs?
arXiv Abstract Not available or no SE tasks 1,531,643 99.85%
Available, with SE tasks 2,330 0.15%
@@
4.1.2 What are the key findings regarding the correct cataloguing of SE PTMs?
The application of our classification pipeline revealed that a significant number of model cards are duplicated. After identifying SE task matches and removing outliers, we found that only 52.11% of the PTMs were actually unique based on their documentation. This duplication is primarily due to two reasons: either they correspond to different variants of the same model (e.g., optimized or finetuned versions), or users re-upload PTMs under their own namespace without making significant modifications to the card content. These duplications reflect the existence of an ecosystem in which PTMs depend on and interact with each other yang2024ecosystemlargelanguagemodels. For example, we observed that sparsh35/Meta-Llama-3.1-8B-Instruct huggingfaceSparsh35MetaLlama318BInstructHugging has 290 highly similar counterparts, such as triplee/torchtune_8B_full_finet
uned_llama3.1_millfield_241223_meta_before_user_epoch2
huggingfaceTripleetorchtune_8B_full_finetuned_llama31_millfield_241223_meta_before_user_epoch2Hugging. Similarly, the PTM dducnguyen/newdream-sdxl-20 huggingfaceDducnguyennewdreamsdxl20Hugging is associated with 715 near-duplicate PTMs. To mitigate this, we implemented a duplicate detection strategy whereby the earliest published model card is treated as the original, and all others with high textual similarity are labeled as similar variants of that.
In addition, during the pilot manual annotation phase, we detected that some tasks could be interpreted in different contexts. To ensure consistency, we used the LLM with explicit inclusion and exclusion criteria for each SE activity, allowing it to disambiguate cases based on context. Notable examples include:
@@ 4.1.2 What are the key findings regarding the correct cataloguing of SE PTMs?
4.2 How are HF PTMs utilized for supporting SE tasks (RQ2)
To answer RQ2, we analyze how PTMs are applied across different SE activities and tasks. We first examine the coverage of SE tasks to identify which areas of the SDLC are supported by PTMs. Next, we explore PTMs that span multiple SE activities to understand cross-activity reuse, and finally, we investigate the relationships between SE activities and ML tasks to highlight patterns of ML support across the software lifecycle.
4.2.1 What SE activities and tasks are covered by PTMs?
Our taxonomy included 147 SE tasks (Section 3.2.1) across the five SE activities of the SDLC, but only a subset of them was represented among the PTMs identified in HF. After completing all phases of our pipeline, we found that 58 SE tasks were covered in the final validated set of 2,205 SE PTMs. Figure 3 shows the distribution of PTMs across these tasks, with each bar color-coded to reflect its corresponding SE activity.
A clear pattern emerges: code generation and coding are by far the most represented SE tasks. Notably, PTMs designed for code generation outnumber those for general coding support by more than a twofold margin. This highlights the strong focus of text-to-code and code-to-code applications of PTMs.
@@ 4.2.1 What SE activities and tasks are covered by PTMs?
From an activity-level perspective, the data reveal that software implementation is the most dominant SE activity in terms of PTM coverage. In contrast, early-stage activities such as requirements engineering and software design are less represented, with only a small fraction of PTMs targeting them. Later phases of the SDLC, including software quality assurance and maintenance, demonstrate moderate PTM adoption, with 150 and 159 PTMs supporting them, respectively. Tasks supported in these phases, such as cybersecurity and sentiment analysis are relatively well-supported, indicating a recognition of the value of automation in ensuring software security.
Figure 3: PTMs associated with each SE task and SE activity
@@
Finding 2.1: Code generation and coding are the most covered SE tasks. Finding 2.2: Software implementation-related PTMs outnumber those for early and late SDLC stages.
4.2.2 How are cross-activity PTMs applied across multiple SE activities?